OpenStudy (anonymous):

could you explain to me what implicit differentiation is? it's pretty annoying.

OpenStudy (anonymous):

post some problems

OpenStudy (anonymous):

yea.. its easier to explain with an example!

OpenStudy (anonymous):

sorry. i have some examples here. x^2 - xy^2 +x^2 +y^2 = 0

OpenStudy (anonymous):

its chain rules

OpenStudy (anonymous):

pretty much, all the other rules still apply , product, quotients etc

OpenStudy (anonymous):

hres the deal trt y and x as seprate functions

OpenStudy (anonymous):

also I think you made a mistake with the question, is there really 2 x^2 terms

OpenStudy (anonymous):

f = g(y) df/dx = g'(y) dy/dx

OpenStudy (anonymous):

its kinda hard to explain over the net, but yeh

OpenStudy (anonymous):

differentiate y with respect to x , you get dy/dx ( obviously )

OpenStudy (anonymous):

Say u have y^5 -y = x, then it's difficult to make y the subject (isolate y, if that's the jargon) and differentiate normally (explicitly). Instead you can 5y^4 dy/dx - dy/dx = 1 and then factor out dy/dx (which is what you are looking for).